OpenStudy (anonymous):

Find f(5) for this sequence: f(1) = 2 and f(2) = 3, f(n) = f(1) + f(2) + f(n - 1), for n > 2. f(5) = ______

OpenStudy (anonymous):

f(n) = f(1) + f(2) + f(n - 1) put n=3 we find f(3) = f(1) + f(2) + f(3 - 1) i.e. f(3) = f(1) + f(2) + f(2) i.e. f(3) = 2 + 2*3 = 8 put n=4 we find f(4) = f(1) + f(2) + f(4 - 1) i.e. f(4) = f(1) + f(2) + f(3) i.e. f(4) = 2 + 3 + 8=13 put n=5 we find f(5) = f(1) + f(2) + f(5 - 1) i.e. f(5) = f(1) + f(2) + f(4) i.e. f(5) = 2 + 3 + 13=18 Thus f(5) = 18 is the required Answer.

OpenStudy (anonymous):

Thank you so much, can you answer another one like it?

OpenStudy (anonymous):

f(1) = 2 and f(2) = 3, f(n) = f(n - 1) + f(n - 2), for n > 2.

OpenStudy (anonymous):

@greenlegodude57 What do you want to find?